calculate the boiling point of a solution containing 12g glucose C6H12O6 dissolved in 200g of water H2O.(0.512oC-K6 of H2O)


1
Expert's answer
2021-06-01T02:45:16-0400

v(c6h12o6)=12g/180g/mol=0.067mol

v(H2O)=200g/18g/mol=11.11mol

t=11.11*0.067/0.512=1.45C